OpenStudy (anonymous):

@misty1212 Convert the following fraction to a decimal. 7/20 A. 0.25 B. 0.35 C. 0.72 D. 0.215

OpenStudy (misty1212):

\[\frac{7}{20}=\frac{7\times 5}{20\times 5}=\frac{35}{100}\] now it is easy to convert

OpenStudy (misty1212):

or you could use a calculator and divide \(7\div 20\) which will also give \(0.35\)

OpenStudy (anonymous):

thanks